a photon of wavelength 12.7 xi 5 m-1 is eminded when it jump from higher to n= 1, what will be higher orbit

a)n=5
b)n=2
c)n=3
d)n=4
1
Expert's answer
2021-09-27T14:13:26-0400

"\frac 1{\lambda}=R(\frac 1{1^2}-\frac{1}{n^2}),\implies"

"n=\frac 1{\sqrt{1-\frac1{\lambda R}}}=2" (ans. b).